# The physical basis of teleological evolutions

### v12

@@ 3.2 Decohered teleological evolutions: what would they look like?
Decohered and endowed with quantum speedup, it is natural to wonder what the supposedly teleological evolutions of life would look like. All natural evolutions are necessarily described in a probabilistic way. The stochastic description of a Darwinian evolution would then be an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical description, transition probabilities can depend on the past of the process, not on the future of it. Instead, a teleological process must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection because more fit to the environment should have the transition toward them more probable than in the classical case.
Let us mention what would be an emblematic example of this kind of natural evolution:
It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
@@
Of course, this teleological explanation is puzzling. However, let us note that it is no more puzzling than the causal loop explanation of the quantum computational speedup. We are in the presence of a quantum cosmological assumption that can make the evolutions of life as puzzling as the microscopic quantum world.
@@ 3.3 How to model teleological natural evolutions
We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth. The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. This obviously holds only in principle, it would however be interesting to investigate the situation by means of toy models.
@@
Let us note that the evolution of life on Earth should have been almost completely independent of the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact that it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[22\right]$.
@@ 3.4 Impact on the natural sciences
The current position of the natural sciences toward teleology can be considered in some measure conflicting. On the one hand, the teleological view would seem to allow a deeper understanding of the functional structures and the evolutions of the living. On the other, as this vision is considered without physical basis and therefore misleading, there is outright hostility to it.
This tension is captured by the following sentence of the evolutionary biologist Haldane $\left[23\right]$:
@@
It would seem that, today, only a few scientists support the teleological view. An example is the biologist Youvan $\left[24\right]$, who shows the revolutionary character that a retrocausal physics would have on the theory of cognition. Another example is that of the philosopher Nagel $\left[25\right]$, who argues that the principles that explain the emergence of life should be teleological rather than materialistic or mechanistic in character.
@@
In recent years, a group of evolutionary biologists $\left[26\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a possible explanation.
As already mentioned, the teleological view would also go along with the fact that some physicists have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[22\right]$.
@@ 3.5 Impact on philosophy
Calling philosophy into play is unconventional in a work of physics. However, we are of course in a very particular case. Indeed, the present quantum theory of teleological evolutions turns out to be the formalization of a notion of natural philosophy dating back two and a half millennia. $\left[27\right]$. We should thus be entitled to look for other possible cases where philosophical thinking has anticipated aspects of the present scientific notion.
@@
For short, let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view777 We can say that what generates the quantum superposition of the time-symmetric causal loops is the mutual causality between the initial and final measurement outcomes. and the causal view. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual causation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[28\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[9\right]$ is a mutual causality view.
@@
Regarding reductionism vs. monism, note that deriving the teleological character of the decohered evolutions of the living on Earth from that of an assumedly quantum evolution of the universe toward life is an essentially monistic view. It presupposes that the description of what we see around and within us is essentially inseparable from that of the whole.
@@
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[29\right]$; in modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. In the present teleological view, this might correspond to the identity and mutual causality between the fundamental physical laws (and their fundamental constants) since the beginning of time and the ideas of our mind. This should also go along with Fritjov Capra’s $\left[30\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics. By the way, we do not need to be anthropocentric. The above would also apply to spatial intelligence, often superior to ours in non-human animals.
@@
(ii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will produce in the future, could also be seen in terms of a “systematically lucky” evolution, systematically lucky in reaching that goal because attracted by it. Interestingly, also the notion of a systematically lucky evolution might have a precedent in pre-scientific thinking. It could be the Stoic notion of $\pi\rho o\nu o\iota\alpha$ (foreknowledge)$\left[31\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge. Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does with quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
(iii) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[32\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today?
@@ 4 Research prospects
(i) Look for a possible synthesis between the present notion of the time-symmetric causal loops and time-symmetric quantum mechanics $\left[33-37\right]$ to which it is obviously inspired.
(ii) The advanced knowledge rule provides the number of oracle queries needed to solve in an optimal quantum way any oracle problem. The unifying character of this rule might suggest the possibility that there is a universal optimal quantum algorithm for this class of problems. Looking for this possibility could be another research prospect.
@@
(vi) Look for applications of the teleological view to quantum self-organization. For example, one could look for toy models of the leap from a soup of life precursor molecules to a self-replicating system, a possibility considered in abiogenesis. Seeing this awkward (classical) leap as the part of a unitary evolution with an attractor in a future full fledged self-replicating system might be appealing.
@@
The above prospects are imagineable extensions of the present work. However, the notion that the quantum framework can host a special form of retrocausality might have a broader impact.
@@ 5 Positioning the work
Of course, we are not alone in challenging the predominant idea that causality always goes in the same (forward) time-direction. Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[33-37\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[36\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[38-42\right]$.
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the approach. In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[39,40\right]$. For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation. The two approaches seem to address different kinds of causality violations.
@@
One is the celebrated EPR paper $\left[3\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description of quantum nonlocality is incomplete. As we have seen, the causal loops generated by time-symmetrizing the ordinary quantum description of quantum nonlocality play the role of the hidden variables they envisaged to physically (locally) explain it $\left[8\right]$.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[43\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separates the two subsystems, until these are not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
@@
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that – for the first time in the history of physics – resorted to retrocausality – see $\left[44\right]$ for an account of this. This tells us a lot about the position of physicists, even at the top level, toward retrocausality. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[28,45\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[9\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a unitary evolution of the universe in between. Dating to before Deutsch’s 1985 discovery of the quantum speedup, of course the two principles in question make no reference to it.
@@ 6 Conclusions
First, we should adopt an ordinary quantum description of quantum algorithms consisting of an initial measurement of the problem-setting in a quantum superposition or mixture of all the possible problem-settings, the unitary computation of the corresponding solution, and the final measurement of the latter.
Then, we time-symmetrize this ordinary quantum description by requiring that the determination of the problem-setting, and thus also that of the corresponding solution, evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. For each possible way of evenly sharing, the part of the selection ascribed to the initial measurement unitarily propagates forward in time until becoming the state immediately before the final measurement; that performed by the final measurement propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome. We should then take the quantum superposition of all these time-symmetrization instances.
@@
This quantum causal loop description of quantum algorithms immediately generalizes to a most general form of quantum correlation, that is, any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms). The ordinary quantum description of any such process is incomplete and is completed by time-symmetrizing it.
@@
We have seen that a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle shows that the evolution of the universe under the fundamental physical laws and constants builds up a quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today (in particular, “existence of sentient life” or “absence of life”). It can correspondingly be endowed with quantum speedup and consequently be teleological in character. As we have seen, in the universe we are in, this possibility would be inherited by: (i) the evolution of the universe toward life, (ii) even if decohered, the evolution of life on Earth, and (iii) the evolutions of species. Under the conservative assumption that the search for the value of life by the evolution of the universe under the fundamental physical law and constants is random, this possibility of quantum speedup becomes an actual quantum speedup.
@@
Explanations of physical phenomena involving retrocausality have been proposed in the past, in particular by Wheeler and Feynman $\left[46\right]$, Costa de Beauregard $\left[43\right]$, Aharonov, Cohen, and Elitzur $\left[36\right]$, and Cramer $\left[47\right]$. Intriguing as they are, it can be said that these explanations never took off, as no form of retrocausality ever entered mainstream quantum mechanics.
The present work makes a similar attempt in the new context of the quantum computational speedup. This should make a difference: the key fact is that any quantum speedup with respect to what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops whose quantum superposition must be implicit in the quantum algorithm that delivers it. This also highlights the teleological character of quantum algorithms, their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (again the solution).

### v13

@@ 4.2 Decohered teleological evolutions: what would they look like?
Decohered and endowed with quantum speedup, it is natural to wonder what the supposedly teleological evolutions of life would look like. All natural evolutions are necessarily described in a probabilistic way. For example, the stochastic description of a Darwinian evolution can be an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical description, the transition probabilities can depend on the past of the process, not on the future of it. Instead, a teleological process must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection because more fit to the environment should have the transition toward them more probable than in the classical case.
Let us mention what could be an emblematic example of this kind of natural evolution:
It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight many millions of years in advance?
@@
Of course, this teleological explanation of the evolution of feathers is puzzling. However, let us note that it is no more puzzling than the causal loop explanation of the quantum computational speedup. We are in the presence of a quantum cosmological assumption that can make the evolutions of life as puzzling as the phenomena of the microscopic quantum world.
@@ 4.3 How to model teleological natural evolutions
We address the problem of how to model a decohered evolution endowed with quantum speedup – as that of the living on Earth should be. Of course, the canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. Although this, in the present case, obviously holds only in principle, it would be interesting to investigate toy models of such an evolution.
@@
Let us note that the evolution of life on Earth should have been almost completely independent of the evolution of the rest of the universe. If it were completely independent, under the quantum cosmological assumption it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact that it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[23\right]$.
@@ 4.4 Impact on the natural sciences
The current position of the natural sciences toward teleology can be considered in some measure conflicting. On the one hand, the teleological view would seem to allow a deeper understanding of the functional structures and the evolutions of the living. On the other, as this vision is considered without any possible physical basis and therefore misleading, there is outright hostility to it.
This tension is captured by the following sentence of the evolutionary biologist Haldane $\left[24\right]$:
@@
It would seem that, today, only a few scientists support the teleological view. An example is the biologist Youvan $\left[25\right]$, who shows the revolutionary character that a retrocausal physics would have on the theory of cognition. Another example is that of the philosopher Nagel $\left[26\right]$, who argues that the principles that explain the emergence of life should be teleological rather than materialistic or mechanistic in character.
@@
In recent years, a group of evolutionary biologists $\left[27\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a possible explanation.
As already mentioned, the teleological view would also go along with the fact that some physicists have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[23\right]$.
@@ 4.5 Impact on philosophy
Calling philosophy into play is unconventional in a work of physics. However, we are of course in a very particular case. Indeed, the present quantum theory of teleological evolutions turns out to be the formalization of a notion of natural philosophy dating back two and a half millennia. $\left[28\right]$. We should thus be entitled to look for other possible cases where philosophical thinking has anticipated aspects of the present scientific notion.
@@
For short, let us call the present teleological view of evolutions and the causal view of today’s physics and the natural sciences respectively the mutual causality view666 We can say that what generates the quantum superposition of the time-symmetric causal loops is the mutual causality between the initial and final measurement outcomes. and the causal view. The causal view would underlie the materialist, mechanistic and reductionist views, while the mutual causality view would underlie the idealist, teleological and monist views.
Consider the opposition between materialism and idealism. The former, about the observer being causally created by physical reality, is a causal view; the latter, about the mutual causation between observer and physical reality, like in the participatory Anthropic Principle of Wheeler $\left[29\right]$, and the anthropic cosmological principles of Barrow and Tipler $\left[8\right]$ is a mutual causality view.
@@
Regarding reductionism vs. monism, note that deriving the teleological character of the decohered evolutions of the living on Earth from that of an assumed quantum evolution of the universe toward life is an essentially monistic view. It presupposes that the description of what we see around and within us is essentially inseparable from that of the whole.
@@
(i) For Plato, “Ideas [of our mind] are objective perfections that exist in themselves and for themselves and, at the same time, they are the causes of natural phenomena and constitute their unity” $\left[30\right]$; in modern language, the cause of natural phenomena that constitute their unity are the fundamental physical laws. In the present teleological view, this might correspond to the identity and mutual causality between the fundamental physical laws (and their fundamental constants) since the beginning of time and the ideas of our mind. This should also go along with Fritjov Capra’s $\left[31\right]$ notion that there is identity between the fundamental states of consciousness described by eastern theosophies and the fundamental laws of quantum physics. By the way, we do not need to be anthropocentric. The above would also apply to spatial intelligence, often superior to ours in non-human animals.
@@
(ii) The present notion of teleological evolution, namely of an evolution toward a goal with an attractor in the goal it will produce in the future, could also be seen in terms of a “systematically lucky” evolution, lucky in reaching that goal because attracted by it. Interestingly, also the notion of a systematically lucky evolution seems to have a precedent in pre-scientific thinking. It could be the Stoic notion of $\pi\rho o\nu o\iota\alpha$ (foreknowledge)$\left[32\right]$, sometimes also translated into providence and, in the present context, possibly translatable into advanced knowledge (e. g., of half of the information about the solution of a problem). Foreknowledge, or advanced knowledge, would have pushed the evolution toward its goal in a luckier and thus faster way than what is possible in the case of a causal evolution, exactly as the advanced knowledge about the solution does with quantum algorithms. This notion of $\pi\rho o\nu o\iota\alpha$ would also fit well with the notion of teleological evolution understood as a stochastic processes with transition probabilities time-symmetrically affected by the past and future of the process, with the luckier transitions more probable than in the classical case.
(iii) Michelangelo’s “Creation of Adam” on the ceiling of the Sistine Chapel could be the manifesto of the present theory. There is a mirror symmetry between creator and created $\left[33\right]$. Could this be interpreted as the time-symmetry – and the related mutual causation – between the fundamental physical laws with their physical constants since the beginning of time and the existence of sentient life today?
@@ 5 Research prospects
(i) Look for a possible synthesis between the present notion of the time-symmetric causal loops and time-symmetric quantum mechanics $\left[34-38\right]$ to which it is obviously inspired.
(ii) The advanced knowledge rule provides the number of oracle queries needed to solve in an optimal quantum way any oracle problem. The unifying character of this rule might suggest the possibility that there is a universal optimal quantum algorithm for this class of problems. Looking for this possibility could be a research prospect.
@@
(vi) Look for applications of the teleological view to quantum self-organization. For example, one could look for toy models of the leap from a soup of life precursor molecules to a self-replicating system, a possibility considered in abiogenesis. Seeing this awkward (classical) leap as the part of a unitary quantum evolution with an attractor in a future full fledged self-replicating system might be appealing.
@@
The above prospects are imagineable extensions of the present work. However, the notion that the quantum framework can host a special form of retrocausality might have a broader impact on the physical view of reality.
@@ 6 Positioning the work
Of course, we are not alone in challenging the predominant idea that causality always goes in the same (forward) time-direction. Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[34-38\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[37\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[39-43\right]$.
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the approach. In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[40,41\right]$. For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation. The two approaches seem to address different kinds of causality violations.
@@
One is the celebrated EPR paper $\left[7\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description of quantum nonlocality is incomplete. As we have seen, the causal loops generated by time-symmetrizing the ordinary quantum description of quantum nonlocality play the role of the hidden variables they envisaged to physically (locally) explain it $\left[5\right]$.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[44\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separates the two subsystems, until these are not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
@@
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie’s, who forbade him to publish that work for years. This, until Wheeler and Feynman published their emitter-absorber theory, which - for the first time in the history of physics - made use of retrocausality - see $\left[45\right]$ for an account of this. This tells us a lot about the position of physicists, even at the top level, towards retrocausality. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[29,46\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[8\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a unitary evolution of the universe in between. Dating to before Deutsch’s 1985 discovery of the quantum speedup, of course the two principles in question make no reference to it.
@@ 7 Conclusions
This cannot be done on the usual quantum description of quantum algorithms – a unitary computation of the solution followed by its measurement – which is itself trivially incomplete lacking the initial measurement. First, we must extend it to the upstream process of setting the problem. An initial measurement selects the problem-setting out of a quantum superposition (or mixture) of all the possible problem-settings; the unitary computation and measurement of the corresponding solution follow. We call this extended quantum description the ordinary quantum description of the quantum algorithm, ordinary because is still a causal description that cannot describe causal loops.
Then, we time-symmetrize this ordinary quantum description by requiring that the determination of the problem-setting, and thus also that of the corresponding solution, evenly shares between the initial measurement of the problem-setting and the final measurement of the solution. For each possible way of evenly sharing, the part of the selection ascribed to the initial measurement unitarily propagates forward in time until becoming the state immediately before the final measurement; the part performed by the final measurement propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous outcome of the initial measurement. We should then take the quantum superposition of all these time-symmetrization instances.
@@
This quantum causal loop description of quantum algorithms immediately generalizes to a most general form of quantum correlation, that is, any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (the problem-setting in quantum algorithms). The time-symmetrization of the ordinary quantum description of this kind of process is imposed by the fact the initial measurement can be postponed at will. When made simultaneous with the final measurement, the time-symmetrization in question must already be in place to satisfy the symmetry implicit in the simultaneous measurements of two maximally entangled observables.
@@
As we have seen, a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle shows that the evolution of the universe under the fundamental physical laws and constants builds up a quantum correlation between the value of the fundamental constants since the beginning of time and the “value” of life today (in particular, “existence of sentient life” or “absence of life”). It can correspondingly be endowed with quantum speedup and thus be teleological in character. In the universe we are in, this possibility would be inherited by: (i) the evolution of the universe toward life, (ii) even if decohered, the evolution of life on Earth, and (iii) the single evolutions on Earth, as those of species. Under the conservative assumption that the search for the value of life by the evolution of the universe under the fundamental physical law and constants is random, the possibility of quantum speedup becomes an actual quantum speedup.
@@
Explanations of physical phenomena involving retrocausality have been proposed in the past in particular by Costa de Beauregard $\left[44\right]$, Wheeler and Feynman $\left[10\right]$, Aharonov, Cohen, and Elitzur $\left[37\right]$, and Cramer $\left[48\right]$. Intriguing as they are, it can be said that these explanations never took off, as no form of retrocausality ever entered mainstream quantum mechanics.
The present work makes a similar attempt in the new context of the quantum computational speedup. This makes two significant differences:
(i) One is the fact is that any quantum speedup with respect to what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops whose quantum superposition must be implicit in the quantum algorithm that delivers it. This also highlights the teleological character of quantum algorithms, their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (again the solution).
(ii) The other is that the retrocausality implicit in the teleological notion, mathematically, is always replaceable by the interplay between quantum state superpositions and quantum interferences that characterize unitary evolutions, indeed the ordinary quantum description of the quantum algorithms. This shouldmake the notion of retrocausality much more palateable.